# Understanding Nonlinear Implicit Bias via Region Counts in Input Space

### v2

@@ 3 Motivation
$d(\cdot)$ and $\gamma(\cdot)$ are commonly used indicator for implicit bias of linear models (Soudry et al., 2018; Ji & Telgarsky, 2018b). However, both of them are not invariant to network reparameterization. We can construct a different set of network weight parameters by scaling the parameters as $\hat{W}=\{2W_{1},\frac{1}{2}W_{2},\cdots,W_{l}\}$, such that $f_{W}=f_{\hat{W}}$ but $d(W)\neq d(\hat{W})$ in general. Similarly, we can scale the last layer weights and get $\tilde{W}=\{W_{1},\cdots,2W_{l}\}$, such that $\gamma(\tilde{W})\neq\gamma(W)$, but $\text{argmax}_{i}f_{W}(x)=\text{argmax}_{i}f_{\tilde{W}}(x)$. This reparameterization trick also works for more complicated norm-based and margin-based generalization metric in (Jiang et al., 2019), or the sharpness-based metrics (Andriushchenko et al., 2023).
@@ 5 Region Counts Correlate with Generalization Gaps
In this section, we present our major empirical findings, which reveal a strong correlation between region counts and the generalization error of neural networks.
@@ 7 Ablation Studies
Mixup (Zhang et al., 2017b) is a data augmentation technique that creates training samples by linearly interpolating pairs of input data and their corresponding labels. We train a ResNet-18 model using mixup, with other hyperparameters in Table 1. The plot in Figure 6 illustrates that Mixup induces smoother decision boundaries with smaller region count and has a better generalization performance. Random crop and random horizontal flip is another way to enhance the diversity of the training dataset. We apply random crop of size 32$\times$32 with padding 4 and random horizontal flip with a probability of 0.5 as data augmentations. As depicted in Figure 7, we observe that compared with mixup, random crop and random flip result in a more evident vertical shift in the performance curve.
@@
Evolution of Region Counts during Training.
Next we study how the region count, generalization gap and their correlation evolve during training. Following the setup in Section 5, we train a ResNet18 model on CIFAR-10 dataset, and report the region count and generalization gap during the training process. The statistics are averaged over different hyperparameter choices as in Table 1.
The results are provided in Table 3. We recorded the average values of region count and generalization gap for these data points in the second and third columns to show their changes during the training process. We observe that the correlation is very low at initialization, but steadily increases during training. This suggests that the metric of region count is not a property of the neural network initialization, but rather inherently involved with the neural network training algorithm.
Table 3: The evolution of region count, generalization gap and their correlation. The correlation is very low at initialization, but steadily increases during training.
Epoch Region Counts Generalization Gap Correlation
0 1.13 N/A N/A
20 3.14 11.2 -0.53
40 3.02 7.3 -0.29
60 3.10 18.2 0.35
80 3.22 26.7 0.77
100 3.25 30.4 0.98
130 3.28 31.2 0.97
160 3.27 31.7 0.98
200 3.26 31.6 0.98
@@ Appendix B How to Calculate the Number of Regions
0:  Prediction Matrix $P$ with dimension$(w,h)$.
0:  Number of connected regions $N$.
@@
10$\times$10 2.76
20$\times$20 2.76
30$\times$30 2.78
40$\times$40 2.78
50$\times$50 2.78
@@ Appendix D Proof
Recall that for each $21\leq i\leq R-$ 2, we have $(w_{n(i)}^{\top}x_{a})(w_{n(i)}^{\top}x_{b})<0$. Therefore, there exists $R-2$ neurons that are activated for either $x_{a}$ or $x_{b}$. This gives $N(x_{a},W)+N(x_{b},W)\geq R-2$, which completes the proof.

### v3

@@ 3 Motivation
[…]
@@ 5 Region Counts Correlate with Generalization Gaps
In this section, we present our major empirical findings, which reveal a strong correlation between region counts and the generalization error of neural networks [^1] Code is available at https://github.com/lijingwei0502/implicit_bias..
@@ 7 Ablation Studies
Mixup (Zhang et al., 2017b) is a data augmentation technique that creates training samples by linearly interpolating pairs of input data and their corresponding labels. We train a ResNet-18 model using mixup, with other hyperparameters in Table 1. The plot in Figure 6 illustrates that Mixup induces smoother decision boundaries with smaller region count and has a better generalization performance. Random crop and random horizontal flip is another way to enhance the diversity of the training dataset. We apply random crop of size $\times32$ with padding 4 and random horizontal flip with a probability of 0.5 as data augmentations. As depicted in Figure 7, we observe that compared with mixup, random crop and random flip result in a more evident vertical shift in the performance curve.
@@ Appendix B How to Calculate the Number of Regions
0: Prediction Matrix $P$ with dimension(w,h)(w,h).
0: Number of connected regions NN.
@@
$\times10$ 2.76
$\times20$ 2.76
$\times30$ 2.78
$\times40$ 2.78
$\times50$ 2.78
@@ Appendix D Proof
Recall that for each $1\leq i\leq R-2$, we have $(w_{n(i)}^{\top}x_{a})(w_{n(i)}^{\top}x_{b})<0$. Therefore, there exists R2R-2 neurons that are activated for either $x_{a}$ or $x_{b}$. This gives $N(x_{a},W)+N(x_{b},W)\geq R-2$, which completes the proof.